Determine the margin of safety in dollars if unit sales = 5,000 units; Dollar Sales = $500,000; Fixed Costs = $204,000; and Variable Costs = $187.5001
Answer #1

Margin of Safety in dollars = $173,600

Workings

Margin of Safety in dollars = Budgeted Sales - Break Even Sales

= 500,000 - 326,400 = $173,600

Break Even Sales = Fixed Costs/Contribution Margin = 204000/62.5% = $326,400

Contribution Margin = (Sales - Variable Costs)/Sales =(500000-187500)/500000 = 312500/500000 = 62.5%

Please hit LIKE button if this helped. For any further explanation, please put your query in comment, will get back to you.